Find the value of r so that line through (-4, 8) and (r, -6) has a alope of 2/3


1
Expert's answer
2020-11-02T20:00:53-0500

Slope =(-6-8)/(r--4)

=-14/r+4

Therefore, -14/r+4=2/3

From the equation, -42=2r+8

2r= -42-8

2r=-50

r=-25.